OpenStudy (anonymous):

find x. 2^((x+6)/2)+2^(x+2)=5

OpenStudy (anonymous):

x is -2

OpenStudy (aum):

You can solve this by trial and error and some guess work. 2 is raised to some power and then added to another 2 raised to some power to give an odd value of 5. One possibility is 4 + 1 = 5 which implies x + 2 = 0 or x = -2 And if we put x = -2 in (x+6)/2 we get 2 and so 2^2 + 2^0 = 5

OpenStudy (anonymous):

u can get an exact solution if u use integer method as i got

OpenStudy (raden):

2^((x+6)/2 ) + 2^(x+2) = 5 2^(x/2 + 3) + 2^(x+2) = 5 2^x/2 * 2^3 + 2^x * 2^2 - 5 = 0 4(2^x) + 8(2^x/2) - 5 = 0 (2*2^x/2 -1 ) (2*2^x/2 + 5) = 0 take each factor equals zero, then solve for x

OpenStudy (raden):

2*2^x/2 -1 = 0 2*2^x/2 = 1 2^x/2 = 1/2 2^x/2 = 2^-1 x/2 = -1 x = -2

OpenStudy (anonymous):

\(\Large 2^{\frac{x+6}{2}}+2^{x+2}=5\) \(\Large \Rightarrow 2^{\frac{x+2}{2}}\times 2^{\frac{4}{2}} +2^{2(\frac{x+2}{2})}=5 \) Let \(\large t=2^{\frac{x+2}{2}}\), the equation become: \(\Large 4t+t^{2}=5\\\Large \Rightarrow t^2+4t-5=0\\\Large \Rightarrow t=1\ \ or\ t=-5 \) \(\Large {\color{red}{t=1}} \Rightarrow 2^{\frac{x+2}{2}}=1 \Rightarrow \frac{x+2}{2}=log_{2}{1}\) \(\Large {\color{red}{t=-5}} \Rightarrow 2^{\frac{x+2}{2}}=-5\Rightarrow \frac{x+2}{2} =log_{2}{-5}\)

OpenStudy (aum):

Last line: Can't take logarithm of a negative number.

OpenStudy (anonymous):

@RadEn no, that is 4t because \(\Large \Rightarrow 2^{\frac{x+2}{2}}\times 2^{\frac{4}{2}} =2^2\times2^{\frac{x+2}{2}}=4t\)

OpenStudy (anonymous):

@aum yup

OpenStudy (anonymous):

so \[\log_{2}1=0 \]x+2=0 x=-2

OpenStudy (anonymous):

yeah! great @amirreza1870 :)